# Searching for the Tetraneutron Resonance on the Lattice

### 15 Apr 2026 v2

@@ III Results and discussion
Recent studies of the $4n$ energy distribution using the 8He($p,p^{4}$He)$4n$ reaction model suggest that the sharp low-energy peak observed by Duer:2022ehf can be explained by dineutron-dineutron correlations Lazauskas:2022mvq. In a finite volume, the $2n$ subsystem becomes quasibound in the sense that the lowest $S$-wave level is shifted to negative energy relative to the two-neutron threshold, even though the infinite-volume nnnn system corresponds to a virtual state. This motivates us to study $2n2n$ scattering phase shifts using Lüscher’s finite-volume method, following Ref. Elhatisari:2016hui. In this context, we employ a finite-volume dineutron approximation, which means that the two-neutron subsystem is treated as an effective weakly bound dimer inside the box, characterized by a large scattering length.
@@
Before turning to $2n2n$ scattering, we validate this finite-volume strategy by studying nnnn and nn2n2n scattering. Since at low energies and large separations the interaction details are then suppressed, we consider the zero-range leading interaction here. We tune the coupling constant of the zero-range neutron-neutron interaction so that the two-neutron subsystem is close to the threshold, and we parameterize this proximity by an effective “binding” scale $B_{2n}$ spanning $0.01$ to $0.09$ MeV. We choose this range to explore shallow finite-volume dimers close to the nnnn threshold. The lower end $B_{2n}=0.01$ MeV deliberately probes the extreme near-threshold regime, where finite-volume and lattice-spacing effects are amplified (see discussion below), while the upper end $B_{2n}=0.09$ MeV remains sufficiently close to threshold to stay in the large-scattering-length regime but provides a practical range for systematic checks.
@@
For each box size $L$ we compute the four-neutron energy $E_{4n}^{(L)}$ and the two-neutron energy $E_{2n}^{(L)}$ in the same periodic volume. The latter defines the effective finite-volume “binding” of the quasibound $2n$ level, $B_{2n}^{(L)}\equiv-E_{2n}^{(L)}$. To extract the relative momentum $p$ of the two-dineutron system we invert the finite-volume dispersion relation for composite dineutrons [Eq. (6)]. This inversion requires as input the infinite-volume dineutron binding energy $B_{2n}$. For the N3LO interaction this scale tends to zero as $L\!\to\!\infty$ and cannot be determined directly from our largest volumes. We therefore treat $B_{2n}$ as an auxiliary near-threshold parameter and scan representative values $B_{2n}=0.01$$0.09$ MeV (all below the smallest $B_{2n}^{(L)}$ encountered at our largest $L$), using the resulting spread as an estimate of the residual systematic uncertainty associated with employing the finite-volume dineutron picture when the underlying nnnn pole is virtual.
@@ IV Summary
To interpret possible $4n$ correlations in terms of two-dineutron dynamics, we employ the finite-volume dineutron approximation, where the lowest nnnn $S$-wave level in a periodic box is quasibound even though the infinite-volume nnnn system corresponds to a virtual state. Within this framework, we validate the Lüscher-based extraction by studying nnnn and nn2n2n scattering. The ratio of scattering lengths $a_{n-2n}/a_{n-n}$ exhibits the expected universal behavior for near-threshold systems, with deviations only when $B_{2n}\approx 0.01$ MeV, where the proximity to the virtual-state regime amplifies finite-volume and lattice-spacing effects.
@@ V Acknowledgments
We are grateful for discussions with Kouichi Hagino, Hiroyuki Sagawa, Furong Xu, and the members of the NLEFT Collaboration. This work is supported by National Natural Science Foundation of China under Grant No. 12435007 and Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China. The work of S.E. is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of U.-G. M. was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of Y. K. was supported in part by the Institute for Basic Science (IBS-R031-D1). L.W. gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University, the National Supercomputing Center of Korea with supercomputing resources including technical support (KSC-2024-CHA-0001, KSC-2025-CHA-0004). $S.$ and U.-G. M. gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at the Jülich Supercomputing Centre (JSC).
@@ V.4 Phase shift
To test the universality of the dineutron approximation under pure contact interactions, we varied the dineutron binding energy from 0.1 MeV down to nearly zero at lattice spacings $a=1.32,\;1.64,\;1.97\;\text{fm}$, computed the relation between the dimensionless quantity $ap\cot\delta$ and apap as shown in Fig. S6. Except for the system with a binding energy of 0.01 MeV, all results align with the expectations, indicating the validity of the dineutron approximation.

### 16 Apr 2026 v3

@@ III Results and discussion
Recent studies of the $4n$ energy distribution using the 8He($p,p^{4}$He)$4n$ reaction model suggest that the sharp low-energy peak observed by Duer:2022ehf can be explained by dineutron-dineutron correlations Lazauskas:2022mvq. In a finite volume, the $2n$ subsystem becomes quasibound in the sense that the lowest $S$-wave level is shifted to negative energy relative to the two-neutron threshold, even though the infinite-volume $nn$ system corresponds to a virtual state. This motivates us to study $2n2n$ scattering phase shifts using Lüscher’s finite-volume method, following Ref. Elhatisari:2016hui. In this context, we employ a finite-volume dineutron approximation, which means that the two-neutron subsystem is treated as an effective weakly bound dimer inside the box, characterized by a large scattering length.
@@
Before turning to $2n2n$ scattering, we validate this finite-volume strategy by studying nnnn and nn2n2n scattering. Since at low energies and large separations the interaction details are then suppressed, we consider the zero-range leading interaction here. We tune the coupling constant of the zero-range neutron-neutron interaction so that the two-neutron subsystem is close to the threshold, and we parameterize this proximity by an effective “binding” scale $B_{2n}$ spanning $0.01$ to $0.09$ MeV. We choose this range to explore shallow finite-volume dimers close to the $nn$ threshold. The lower end $B_{2n}=0.01$ MeV deliberately probes the extreme near-threshold regime, where finite-volume and lattice-spacing effects are amplified (see discussion below), while the upper end $B_{2n}=0.09$ MeV remains sufficiently close to threshold to stay in the large-scattering-length regime but provides a practical range for systematic checks.
@@
For each box size $L$ we compute the four-neutron energy $E_{4n}^{(L)}$ and the two-neutron energy $E_{2n}^{(L)}$ in the same periodic volume. The latter defines the effective finite-volume “binding” of the quasibound $2n$ level, $B_{2n}^{(L)}\equiv-E_{2n}^{(L)}$. To extract the relative momentum $p$ of the two-dineutron system we invert the finite-volume dispersion relation for composite dineutrons [Eq. (6)]. This inversion requires as input the infinite-volume dineutron binding energy $B_{2n}$. For the N3LO interaction this scale tends to zero as $L\!\to\!\infty$ and cannot be determined directly from our largest volumes. We therefore treat $B_{2n}$ as an auxiliary near-threshold parameter and scan representative values $B_{2n}=0.01$$0.09$ MeV (all below the smallest $B_{2n}^{(L)}$ encountered at our largest $L$), using the resulting spread as an estimate of the residual systematic uncertainty associated with employing the finite-volume dineutron picture when the underlying $nn$ pole is virtual.
@@ IV Summary
To interpret possible $4n$ correlations in terms of two-dineutron dynamics, we employ the finite-volume dineutron approximation, where the lowest nnnn $S$-wave level in a periodic box is quasibound even though the infinite-volume $nn$ system corresponds to a virtual state. Within this framework, we validate the Lüscher-based extraction by studying nnnn and nn2n2n scattering. The ratio of scattering lengths $a_{n-2n}/a_{n-n}$ exhibits the expected universal behavior for near-threshold systems, with deviations only when $B_{2n}\approx 0.01$ MeV, where the proximity to the virtual-state regime amplifies finite-volume and lattice-spacing effects.
@@ V Acknowledgments
We are grateful for discussions with Kouichi Hagino, Hiroyuki Sagawa, Furong Xu, and the members of the NLEFT Collaboration. This work is supported by National Natural Science Foundation of China under Grant No. 12435007 and Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China. The work of S.E. is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of U.-G. M. was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of Y. K. was supported in part by the Institute for Basic Science (IBS-R031-D1). L.W. gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University, the National Supercomputing Center of Korea with supercomputing resources including technical support (KSC-2024-CHA-0001, KSC-2025-CHA-0004). S.S. and U.-G. M. gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at the Jülich Supercomputing Centre (JSC).
@@ V.4 Phase shift
To test the universality of the dineutron approximation under pure contact interactions, we varied the dineutron binding energy from 0.1 MeV down to nearly zero at lattice spacings $a=1.32,\;1.64,\;1.97\;\text{fm}$, computed the relation between the dimensionless quantity $ap\cot\delta$ and $ap$ as shown in Fig. S6. Except for the system with a binding energy of 0.01 MeV, all results align with the expectations, indicating the validity of the dineutron approximation.